OpenStudy (anonymous):

Claire has received scores of 85, 88, 87, and 80 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 82? (Hint: The average of a list of numbers is their sum divided by the number of numbers in the list.)

OpenStudy (anonymous):

\[\frac{ 85+88+87+80+x }{5 }>=82\]

OpenStudy (anonymous):

does it mean x=70

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

410 >= 340 + x x <= 70

OpenStudy (anonymous):

sorry x >= 70

OpenStudy (anonymous):

x can be 70 or any greater number

OpenStudy (anonymous):

question is what is the minimum, it cannot be more then 1 number :)

OpenStudy (anonymous):

ok